Calculate the average kinetic energy of each nitrogen molecule at 273 degree celsius

Grade:11

1 Answers

Arun
6 years ago
Dear student
 
The average kinetic energy of the molecules is 5.66 × 10?²¹ J.
According to Kinetic Molecular Theory, the average kinetic energy of gas molecules is a function only of temperature.
The formula is
#KE = 3/2kT#
where #T# is the Kelvin temperature and #k# is Boltzmann's constant.
#KE = 3/2 × 1.381 × 10?²³ "J·K?¹" × 273"K"# = 5.66 × 10?²¹ J
For a mole of molecules, the average kinetic energy is
#(5.66 × 10?²¹"J")/(1"molecule") × (6.022 × 10²³"molecules")/(1"mol")# = 3.41 ×10³ J/mol = 3.41 kJ/mol

Regards
Arun (askIITians forum expert)